# Ptychoformer: A Physics-Guided Deep Learning Framework for Ptychographic Imaging

### v1

@@ III-B3 Feature Fusion and Decoding
Let $\mathcal{F}_{\text{Local}}:L^{2}(\mathbb{R}^{2}\times\{1,\ldots,J\})\to\mathcal{H}_{\text{Local}}$ and $\mathcal{F}_{\text{NonLocal}}:L^{2}(\mathbb{R}^{2}\times\{1,\ldots,J\})\to\mathcal{H}_{\text{NonLocal}}$ denote the feature extraction operators from the Local Dependencies Branch and the NonLocal Coherence Branch, respectively. The decoder is defined as a composition of three upsampling operators $\mathcal{D}=\mathcal{D}_{3}\circ\mathcal{D}_{2}\circ\mathcal{D}_{1}$, each progressively increasing the spatial resolution using transposed convolutions. Here, $\circ$ denotes function composition, such that $(f\circ g)(x)=f(g(x))$ for any compatible functions $f$ and $g$. The feature fusion process, combining features from both branches, is formulated as:
$\mathcal{F}_{\text{fused}}=\mathcal{C}(\mathcal{F}_{\text{Local}},\mathcal{F}_{\text{NonLocal}})$ (12)
where $\mathcal{C}$ concatenates $\mathcal{F}_{\text{Local}}$ and $\mathcal{F}_{\text{NonLocal}}$, followed by a convolution to integrate the features. The fused features $\mathcal{F}_{\text{fused}}$ are then processed through the three-layer decoder $\mathcal{D}$, implemented using transposed convolutions with batch normalization and ReLU activations, split into parallel branches $\mathcal{D}_{\text{amplitude}}$ and $\mathcal{D}_{\text{phase}}$, to reconstruct the preliminary amplitude $\mathbf{A}(\mathbf{r})$ and phase $\boldsymbol{\phi}(\mathbf{r})$ images.
III-B4 Output Constraints
To ensure the physical validity of the reconstructed complex-valued object function, we apply constraints to the final outputs. The amplitude $\mathbf{A}_{\text{final}}:\mathbb{R}^{2}\to\mathbb{R}^{+}$ and phase $\boldsymbol{\phi}_{\text{final}}:\mathbb{R}^{2}\to[-\pi,\pi]$ are constrained as:
$\mathbf{A}_{\text{final}}(\mathbf{r})=A_{\text{max}}\cdot\sigma\left(\frac{\mathbf{A}(\mathbf{r})}{A_{\text{max}}}\right),\quad\boldsymbol{\phi}_{\text{final}}(\mathbf{r})=\pi\cdot\tanh\left(\boldsymbol{\phi}(\mathbf{r})\right)$ (13)
where $A_{\text{max}}$ is the maximum allowable amplitude, $\sigma$ is the sigmoid function, and $\tanh$ is the hyperbolic tangent function. The final complex-valued object function is reconstructed as
$O_{\text{reconstructed}}(\mathbf{r})=\mathbf{A}_{\text{final}}(\mathbf{r})\cdot\exp\left(i\boldsymbol{\phi}_{\text{final}}(\mathbf{r})\right).$
III-B5 Training Objective
@@
where $N$ is the number of training samples, $A_{\text{true}}^{(n)}$ and $\phi_{\text{true}}^{(n)}$ are the true amplitude and phase for the $n$-th sample, respectively, and $A_{\text{pred}}^{(n)}=\mathbf{A}_{\text{final}}^{(n)}$ and $\phi_{\text{pred}}^{(n)}=\boldsymbol{\phi}_{\text{final}}^{(n)}$ are the corresponding model predictions after applying the output constraints. The optimization problem is formulated as:
@@ IV-A1 Implementation Details
The model was trained using a batch size of 32 for 25 epochs. Adam optimizer with ReduceLROnPlateau callback (factor 0.5, patience 2, $\min_{\operatorname{lr}}$ 0.0001) adjusts learning rate based on validation loss. Training on Tesla T4 GPU. Custom callback monitors memory, time, and convergence. Model compiled with MSE loss. ReduceLROnPlateau strategy implemented. Dataset shuffled with 5% validation split.
@@ IV-A2 Datasets
We utilize two types of datasets in our experiments: Simulated Dataset: Dataset based on magnetic nanostructure simulation [29] using Ni80Fe20 parameters. 512×512 LTEM images represent 5×5 nm areas. The dataset contains 2982 images with 75% overlap and 3-pixel jitter. Gaussian noise simulates realism, with added Poisson and Gaussian noise for readout and photon effects. Amplitude and phase undergo minmax normalization. Real Experimental Data: Experimental data from Argonne National Laboratory [9] comprises 16,100 triplets of diffraction data, amplitude, and phase images. 161 × 161 point scan at 30 nm increments from X-ray nanoprobe beamline 26-ID. PIE algorithm generates ground truth. For both datasets, we split the data into 80% for training and 20% for testing. The validation set is created using 5% of the training data.
@@ IV-A3 Evaluation Metrics
Figure 4: Performance comparison of single-shot experiment results and full-stitched scene retrieval. (a) and (b) show two representative scan points. (c) represents the full-stitched scene retrieval across models.
TABLE I: Quantitative comparison of full-stitched scenes under noise-free and noisy conditions. Results are presented as mean $\pm$ standard deviation from ten independent experiments.
@@
Noise-free PtychoNN 4.32 $\pm$ 0.02 13.52 $\pm$ 0.17 83.50 $\pm$ 0.80 1.22 $\pm$ 0.01 19.09 $\pm$ 0.19 67.10 $\pm$ 0.70
PtyNet 4.46 $\pm$ 0.03 13.56 $\pm$ 0.15 83.10 $\pm$ 0.60 1.28 $\pm$ 0.01 19.10 $\pm$ 0.26 67.80 $\pm$ 1.30
DPI 4.29 $\pm$ 0.03 13.68 $\pm$ 0.12 83.40 $\pm$ 0.50 1.26 $\pm$ 0.01 18.91 $\pm$ 0.24 67.70 $\pm$ 1.10
Ptychoformer 3.63 $\pm$ 0.04 14.42 $\pm$ 0.08 87.00 $\pm$ 0.30 1.14 $\pm$ 0.01 19.41 $\pm$ 0.05 70.60 $\pm$ 0.70
@@ IV-B Performance Comparison and Analysis
We evaluated Ptychoformer against three mainstreamed CNN-based methods (PtychoNN, PtyNet, and DPI). Our analysis covered single-shot retrieval and full-stitched field retrieval under both ideal and noisy conditions.
@@ IV-B1 Single Scan Point Retrieval Analysis
Ptychoformer produced retrievals with improved edge definition and clarity compared to those from the other methods, particularly evident in the vertical edge on the left side of Fig. 4(a). The reconstructed images also showed contrast levels closer to the ground truth, with more natural transitions between dark and bright areas. By contrast, the CNN-based models tended to smooth these sharp features and introducing artifacts and distortions.
@@ IV-B2 Full-stitched Field Retrieval Analysis
Table I presents full-field stitching results under ideal (noise-free) and realistic (noisy) conditions, demonstrating Ptychoformer’s superior performance across all metrics. In ideal conditions, Ptychoformer achieved 16.0% lower MSEamp and 6.7% higher PSNRamp compared to PtychoNN, with similar improvements under noisy conditions. Fig. 4(c) visually confirms Ptychoformer’s superior global consistency and minimal boundary artifacts, particularly in reconstructing vertical column ends and background textures. The dual-branch structure enables our model to effectively capture both low-frequency and mid-to-high frequency information of diffraction patterns. This capability stems from the model’s ability to capture the intrinsic geometric structure of the data, analogous to continuous mapping on high-dimensional manifolds.
Repeated measures ANOVA and pairwise $t$-tests (Bonferroni correction, $\alpha=0.0083$) showed Ptychoformer significantly outperformed the other methods in all metrics in both conditions ($p<0.0083$), especially in MSEamp and SSIMamp ($p<0.0001$). The three CNN-based methods showed no significant differences, performing similarly across most metrics.
@@ IV-C Validation on Experimental Synchrotron Data
Figure 5: Performance comparison on real experimental samples. (a,b) Visual comparison of retrieved amplitude and phase. For a specific hook-shaped detail, only our model effectively restores it, significantly outperforming CNN-based methods in fine structure retrieval. (c-e) Quantitative comparison of PSNR, SSIM, and MSE.
Fig. 5 demonstrates Ptychoformer’s superior performance on real experimental data, consistent with our findings from simulated datasets. Across all metrics (PSNR, SSIM, and MSE) for both amplitude and phase retrieval, Ptychoformer consistently outperforms other models. These improvements are visible in fine structural details (Fig. 5a,b), indicating Ptychoformer’s enhanced resolution and fidelity in practical imaging applications.
@@ IV-D Generalization Capabilities Comparison
We evaluated Ptychoformer’s generalization by trained on simulated tungsten patterns with straight-line features and evaluated on complex, curved patterns without fine-tuning, assessing the model’s ability to extrapolate learned features to significantly different sample geometries. Fig. 6 shows Ptychoformer’s superior retrieval of curved patterns. Quantitative results in Table II support these visual observations. For amplitude retrieval, Ptychoformer achieves a PSNRamp of 8.92 dB and SSIMamp of 59.06%, outperforming other models by 0.83-0.95 dB and 8.51-11.28%, respectively. For phase retrieval, Ptychoformer attains a PSNRphase of 8.68 dB and SSIMphase of 49.44%, surpassing other models by 0.91-1.04 dB and 32.16-35.71%, respectively. These results demonstrate Ptychoformer’s ability to capture fundamental relationships between diffraction patterns and sample structures, aligning with the geometry-agnostic nature of ptychographic phase retrieval.
Figure 6: Comparison of generalization capabilities across different models: Direct test on a new unseen sample with significantly different distribution from the training set. (a) Amplitude and (b) phase retrieval of a single diffraction pattern. (c) Amplitude and (d) phase retrieval of a full scene (Fudan University logo). Models compared: Ground truth, PtychoNN, PtyNet, DPI, and Ptychoformer.
TABLE II: Performance Comparison for Out-of-Distribution Generalization.
Model Amplitude Phase
PSNR (dB) SSIM (%) PSNR (dB) SSIM (%)
PtychoNN 8.03 50.55 7.77 37.34
PtyNet 7.97 47.78 7.64 36.43
DPI 8.09 50.14 7.65 37.41
Ptychoformer 8.92 59.06 8.68 49.44
@@ IV-E Performance Comparison Under Limited Data
Despite significant advancements in ptychography, a major challenge remains: extended data acquisition periods. While higher overlap between scanning points typically yields better retrieval results [14], it also increases experiment duration and radiation exposure. This is particularly problematic for radiation-sensitive materials [30] and in situ dynamic studies [31]. Our study focuses on this aspect, evaluating model performance under two kinds of limited data conditions.
IV-E1 Impact of Training Set Size
Figure 7: Training and testing set divisions of different sizes. Black boxes represent training set sizes, while white boxes represent testing sets, shown as squares.
Fig. 8 (a-c) illustrates Ptychoformer’s performance under varying training set sizes. As the training size decreases from 2928 to 50, Ptychoformer consistently outperforms CNN-based models, especially at smaller training set sizes. Comparing with DPI, the best-performing CNN method, we showcase results at representative training sizes of 1000 and 50. Even with 1000 samples, DPI exhibits limitations with complex structures like the top-right crosshatch pattern, producing block-shaped high-frequency artifacts. Ptychoformer, in contrast, maintains superior resolution. At the minimal training size of 50 samples, DPI’s retrieval quality significantly deteriorates, while Ptychoformer still distinguishes basic structural features like crosshatches and L-shaped patterns. This robustness is attributed to Ptychoformer’s PCA mechanism and dual-branch architecture, which align with diffraction physics and capture multi-scale features effectively, demonstrating superior ability to learn from limited data.
IV-E2 Performance Across Overlap Ratios
Fig. 8 (d) demonstrates Ptychoformer’s effectiveness at varying overlap ratios. At 0% overlap, Ptychoformer achieves higher PSNRamp, surpassing CNNs. From 75% to 0% overlap, Ptychoformer’s PSNRamp drops by a smaller amount compared to CNNs. This resilience, analyzable through compressed sensing theory. [32], suggests Ptychoformer learns an optimal sparse prior. Lower overlap requirements could enable faster experiments with reduced radiation damage. Ptychoformer’s efficiency demonstrates the synergy between deep learning and domain-specific knowledge in ptychography.
Figure 8: Performance comparison across training set sizes and overlap ratios. (a,b) Retrievals with 1000 and 50 training samples. (c,d) PSNR and SSIM metrics for varying training sizes and overlap ratios. Higher values for both PSNR and SSIM indicate better performance.
TABLE III: Ablation Study Results

### v2

@@ III-B3 Feature Fusion and Decoding
The feature fusion process addresses the resolution mismatch between the two branches. The Local Dependencies Branch output $\mathcal{H}_{\text{Local}}\in\mathbb{R}^{B\times(H/P)\times(W/P)\times D}$ has lower spatial resolution than the Non-Local Branch output $\mathcal{H}_{\text{NonLocal}}\in\mathbb{R}^{B\times H\times W\times C^{\prime}}$. To align these representations, we employ bilinear upsampling on the Local Branch features:
$\mathcal{H}_{\text{Local}}^{\text{upsampled}}=\text{Upsample}_{\text{bilinear}}(\mathcal{H}_{\text{Local}})\in\mathbb{R}^{B\times H\times W\times D}$ (9)
The aligned features are then combined through concatenation followed by a 1×1 convolution:
$\mathcal{H}_{\text{fused}}=\text{Conv}_{1\times 1}(\text{Concat}[\mathcal{H}_{\text{Local}}^{\text{upsampled}},\mathcal{H}_{\text{NonLocal}}])$ (10)
The decoder $\mathcal{D}$ consists of three transposed convolution blocks with progressive upsampling:
$\begin{split}\mathcal{D}_{i}&=\text{BatchNorm}(\text{ReLU}(\text{ConvTranspose2D}_{k,s}))\\ \mathcal{D}&=\mathcal{D}_{3}\circ\mathcal{D}_{2}\circ\mathcal{D}_{1}\end{split}$ (11)
where $k$ is kernel size and $s$ is stride. The decoder branches into parallel paths $\mathcal{D}_{\text{amplitude}}$ and $\mathcal{D}_{\text{phase}}$ to reconstruct the object’s amplitude and phase components.
III-B4 Training Objective
@@
where $N$ is the number of training samples, $A_{\text{true}}^{(n)}$ and $\phi_{\text{true}}^{(n)}$ are the true amplitude and phase for the $n$-th sample (both amplitude and phase values are normalized to the range [0,1]), respectively, and $A_{\text{pred}}^{(n)}=A_{\text{final}}^{(n)}$ and $\phi_{\text{pred}}^{(n)}=\boldsymbol{\phi}_{\text{final}}^{(n)}$ are the corresponding model predictions. The two error terms naturally maintain comparable scales without requiring additional weighting factors.The optimization problem is formulated as:
@@ IV-A Experiment Setup
Figure 3: Performance comparison of single-shot experiment results and full-stitched scene retrieval using simulated data. (a) and (b) show amplitude and phase reconstructions at two representative scan positions. (c) displays the full-field amplitude and phase images stitched together from individual reconstructions at all scan positions across different models.
@@ IV-A1 Implementation Details
We implemented both our DL-based method and the traditional iterative algorithm (like the extended ptychographic iterative engine (ePIE) algorithm[29]) in Python with CUDA acceleration on the same hardware platform. The iterative algorithm was implemented with CUDA support, initialized with known probe measurements and processing 32$\times$32 diffraction patterns, with phase unwrapping applied to retrieve the continuous phase. The ePIE implementation typically required 100-200 iterations to converge and 75-80% spatial overlap, with minimal gains beyond. Both implementations were trained and evaluated on a NVIDIA® Tesla T4 GPU with Intel® Xeon® Platinum 8480CL CPU @ 2.00GHz running Linux 6.1.85+. For the DL-based method, we used TensorFlow with an MSE loss function and the Adam optimizer. We set the initial learning rate to 1.0 and implemented a ReduceLROnPlateau callback (factor 0.5, patience 2, $\min_{\operatorname{lr}}$ 0.0001) to dynamically adjust the learning rate based on validation loss. We used a batch size of 32 with 5% validation split, while a custom callback monitored system memory usage, computation time, and convergence behavior.
@@ IV-A2 Datasets
We utilize two types of datasets in our experiments: Simulated Dataset: The dataset consists of a 512×512 LTEM image[30]. The pattern is based on a slightly modified version of the 1951 USAF resolution test chart. Using a probe size of 32×32 pixels with a 75% overlap rate and 3-pixel jitter, we generated a total of 3721 patches arranged in a 61×61 grid, each patch being 32×32 pixels. Gaussian noise simulates realism, with added Poisson and Gaussian noise for readout and photon effects. Amplitude and phase undergo minmax normalization. Real Experimental Data: Experimental data from Argonne National Laboratory [9] comprises 16,100 triplets of diffraction data, amplitude, and phase images. 161 × 161 point scan at 30 nm increments from X-ray nanoprobe beamline 26-ID. PIE algorithm generates ground truth. For both datasets, we split the data into 80% for training and 20% for testing. The validation set is created using 5% of the training data. Results on simulated data are presented in Sections IV.B, D, E, F, G and Section V.B, while Section IV.C validates the model using real experimental synchrotron data.
@@ IV-A3 Evaluation Metrics
TABLE I: Quantitative comparison of full-stitched scenes under noise-free and noisy conditions on simulated data. Results are presented as mean $\pm$ standard deviation from ten independent experiments.
@@
Noise-free PtychoNN[9] 4.32 $\pm$ 0.02 13.52 $\pm$ 0.17 83.50 $\pm$ 0.80 1.22 $\pm$ 0.01 19.09 $\pm$ 0.19 67.10 $\pm$ 0.70
PtyNet[11] 4.46 $\pm$ 0.03 13.56 $\pm$ 0.15 83.10 $\pm$ 0.60 1.28 $\pm$ 0.01 19.10 $\pm$ 0.26 67.80 $\pm$ 1.30
DPI[10] 4.29 $\pm$ 0.03 13.68 $\pm$ 0.12 83.40 $\pm$ 0.50 1.26 $\pm$ 0.01 18.91 $\pm$ 0.24 67.70 $\pm$ 1.10
PPN (Ours) 3.63 $\pm$ 0.04 14.42 $\pm$ 0.08 87.00 $\pm$ 0.30 1.14 $\pm$ 0.01 19.41 $\pm$ 0.05 70.60 $\pm$ 0.70
@@ IV-B Performance Comparison and Analysis
We evaluated PPN against three mainstreamed CNN-based methods (PtychoNN, PtyNet, and DPI). Our analysis covered single-shot retrieval and full-stitched field retrieval under both ideal and noisy conditions.
@@ IV-B1 Single Scan Point Retrieval Analysis
PPN reconstructed retrievals with improved edge definition and clarity compared to those from the other methods in single scan point recovery, where each position represents an individual amplitude/phase reconstruction from a corresponding diffraction pattern. This improvement is particularly evident in the vertical edge on the left side of Fig. 3(a). The reconstructed images also showed contrast levels closer to the ground truth, with more natural transitions between dark and bright areas. By contrast, the CNN-based models tended to smooth these sharp features and introducing artifacts and distortions accuracy.
Figure 4: Frequency analysis comparison of full-scene ptychographic retrievals on simulated data. (a) 1D diagonal cross-sections of average 2D PSD from stitched simulations. Red curve: ground truth; Blue curve: models. The red box represents obvious abnormal model retrieval. (b) Quantitative breakdown of energy distribution across low, mid, and high frequency bands. Frequency ranges are defined based on the radial distance from the PSD center, with boundaries at 1/3 and 2/3 of the maximum frequency.
@@ IV-B2 Full-stitched Field Retrieval Analysis
Table I presents full-field stitching results under ideal (noise-free) and realistic (noisy) conditions, demonstrating PPN’s superior performance across all metrics. In ideal conditions, PPN achieved 16.0% lower MSEamp and 6.7% higher PSNRamp compared to PtychoNN, with similar improvements under noisy conditions. Fig. 3(c) visually confirms PPN’s superior global consistency and minimal boundary artifacts, particularly in reconstructing vertical column ends and background textures. The dual-branch structure enables our model to effectively capture both low-frequency and mid-to-high frequency information of diffraction patterns. This capability stems from the model’s ability to capture the intrinsic geometric structure of the data, analogous to continuous mapping on high-dimensional manifolds.
Repeated measures ANOVA and pairwise $t$-tests (Bonferroni correction, $\alpha=0.0083$) showed PPN significantly outperformed the other methods in all metrics in both conditions ($p<0.0083$), especially in MSEamp and SSIMamp ($p<0.0001$). The three CNN-based methods showed no significant differences, performing similarly across most metrics. Fig. 3 focuses on intra-family comparisons among deep learning architectures to isolate the impact of structural variations under consistent training supervision. Comparative evaluations with conventional iterative algorithms (e.g., ePIE) are presented separately in Section IV-E3, where differences in computational efficiency and resilience to varying overlap ratios are quantitatively assessed and visually demonstrated.
IV-B3 Spatial Frequency Fidelity Analysis
We analyze frequency-domain fidelity using 1D diagonal cross-sections of the averaged power spectral density (PSD), defining three bands by normalized radial distance: low (0-1/3 Nyquist), mid (1/3-2/3), and high (2/3-1) frequencies.
[…]
@@ IV-C Validation on Experimental Synchrotron Data
Figure 5: Performance comparison on real experimental samples. (a,b) Visual comparison of retrieved amplitude and phase (scale bar: 200nm). For a specific hook-shaped detail in ROI (region of interest), only our model effectively restores it, significantly outperforming CNN-based methods in fine structure retrieval. (c-e) Quantitative comparison of PSNR, SSIM, and MSE.
Figure 6: Comparison of generalization capabilities across different methods using the Fudan University logo (simulated data) as a test sample with distribution significantly different from the training set. (a) Phase retrieval results with intensity values normalized to $\pi$. (b) Detailed ROI comparison.
Fig. 5 demonstrates PPN’s superior performance on real experimental data, consistent with our findings from simulated datasets. Across all metrics (PSNR, SSIM, and MSE) for both amplitude and phase retrieval, PPN consistently outperforms other models. These improvements are visible in fine structural details (Fig. 5 (a,b)). The model’s ability to preserve edge sharpness and resolve intricate features in experimental data validates its potential for practical applications where accurate reconstruction of complex nanostructures is essential for scientific interpretation.
@@ IV-D Generalization Capabilities Comparison
We evaluated PPN’s generalization by training on simulated 1951 USAF pattern (the same with Section IV-B) with straight-line features and evaluating on complex, curved patterns (Fig. 6), without fine-tuning, assessing the model’s ability to extrapolate learned features to significantly different sample geometries. This evaluation connects directly to practical applications where pre-trained models must process new samples with unknown structures—a scenario where acquiring specific training data is often impractical. In these contexts, generalization capability determines the practical utility of deep learning for ptychographic imaging. Fig. 6 compares phase retrieval results across different methods. While deep learning models cannot match the reconstruction quality of the ePIE algorithm, PPN outperforms all other DL-based methods. The red boxes highlight high-frequency details where PPN preserves spatial information with minimal artifacts compared to other end-to-end methods.
Results show that even with different feature distributions between training and testing data, PPN achieves superior reconstruction quality and better preservation of high-frequency spatial details.
@@ IV-E Performance Comparison Under Limited Data
Despite significant advancements in ptychography, a major challenge remains: extended data acquisition periods. While higher overlap between scanning points typically yields better retrieval results [14], it also increases experiment duration and radiation exposure. This is particularly problematic for radiation-sensitive materials [31] and in situ dynamic studies [32]. Here, the overlap ratio is defined as the physical overlap between adjacent scanning positions in real space.
Figure 7: Model performance evaluation on simulated data. (a) Reconstruction quality versus training data size at overlap ratio 75%. (b) Performance stability under different test set overlap ratios with fixed training set size (100 samples).
IV-E1 Comparison with Different Training Data Size
We evaluated the effect of training data size on model performance by training models with 10 different random initializations while keeping all hyperparameters consistent and testing on the same test set as in Section IV-B. As shown in Fig. 7(a), when the training samples decrease from 2928 to 50, PPN demonstrates superior data efficiency: the PSNR for amplitude reconstruction decreases by only 35%, and phase reconstruction by 15%. In contrast, PtychoNN shows the most significant performance degradation, with PSNR dropping by approximately 70% for both amplitude and phase reconstruction. This resilience, analyzable through compressed sensing theory. [33], suggests PPN learns an optimal sparse prior.
IV-E2 Comparison Across Overlap Ratios with a Fixed Training Size
We investigated the impact of test set overlap ratios (0%, 25%, 50%, and 75%) while maintaining a fixed training set size of 100 samples, consistent with the models used in Section IV-E1. While the test area maintains consistent spatial coverage (matching Fig. 3 (c)’s field of view), the number of test samples decreases with lower overlap ratios due to reduced spatial sampling density. As illustrated in Fig. 7(b), all models exhibit remarkable metric stability across different overlap ratios. Taking PPN as an example, its amplitude reconstruction PSNR remains stable between 9.5–9.6 ($\Delta=0.1$) and SSIM between 0.39–0.40 ($\Delta=0.01$), with one-way ANOVA ($p>0.05$) confirming no significant differences between overlap ratio groups.
This overlap-invariant behavior stems from our training approach, which uses simulated real-space amplitude and phase as ground truth—data that typically requires exceptionally high overlap ratios to obtain in conventional experimental workflows. This characteristic enables an optimal workflow: using high-overlap data for model training while deploying at low overlap during testing. These findings naturally lead to the question of how PPN compares to traditional iterative methods across different overlap ratios, which we explore in the following section.
Figure 8: Comparison between ietrative method and PPN across different overlap ratios. (a) SSIM and reconstruction time comparison. (b) and (c) Reconstructed images using iterative method and our proposed PPN, respectively, at different overlap ratios (30%, 60%, 90%).
Figure 9: Ablation study performed on simulated data showing amplitude and phase reconstructions with 2D images and 1D cross-sectional profiles , comparing ground truth (blue solid lines) with AI inference results (red dashed lines).
IV-E3 Comparison with Iterative Methods Across Overlap Ratios
We compared PPN with the iterative method (where we use the widely-used ePIE as our baseline) across various overlap ratios to assess both reconstruction quality and computational efficiency. Fig. 8 presents this comparison, with PPN trained using the same dataset described in Section IV-B, where sample features remain within the learned distribution. As shown in Fig. 8(a), PPN maintains SSIM values between 0.86–0.92 across all overlap ratios, indicating consistent reconstruction quality. In contrast, iterative method exhibits strong overlap dependence, with SSIM dropping from 0.95 at 90% overlap to only 0.12 at 30% overlap. At 30% overlap, PPN completes reconstructions in approximately 0.15 seconds versus iterative method’s 125 seconds, representing an 852$\times$ speed improvement, and at 60% overlap, a 1767× speed improvement. The reconstruction time was measured by averaging over ten consecutive tests, with error bars representing the variance. Reducing overlap from 90% to 30% increases data acquisition efficiency by approximately 49 times for a fixed area, as scan point density is inversely proportional to the square of the step size. Theoretically, this combined effect could yield up to 41,748$\times$ (49 $\times$ 852) overall efficiency improvement. This enables significantly faster experiments with substantially reduced radiation exposure while still producing high-quality reconstructions.
This separation of training and testing strategies, along with PPN’s efficiency, demonstrates the synergy between deep learning and domain-specific knowledge in ptychography, providing a viable technical pathway for achieving high-throughput imaging.
IV-F Ablation Study
TABLE II: Ablation Study Results on Simulated Data